From 16 litre milk 2 litre milk was taken out and 2 litre of water was added in it.And this process was done 3 times.What will be the ratio of concentration of milk and water in the resulting solution?

Grade:12

2 Answers

rohit rathi
13 years ago

15:28  is the answer.

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Vedanshu

After first process quantity of milk=14 lit and water =2 lit

After second process quantity of milk=14*(1-2/16)

As one can observe easily that concentration of milk is reduced by 1/8 each time

so conc of milk=(7/8)^3

conc of water=1-(7/8)^3

So ratio of two =343/169

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..